### v1

@@ 3.1.3 Placebo effects
$\text{AVSQ}_{-\ell}=E\left[S(Y_{F-1-\ell}-Y_{F-1}-\Delta_{-\ell}(D_{1},F))|\bm{D}\in\mathcal{D}^{\ell},F-1-\ell\geq 1\right],$
@@ 3.1.5 Other estimators and identifying strategies
Our estimator of $\text{AVSQ}_{\ell}$ consists in estimating the missing counterfactual outcome $Y_{g,F_{g}-1+\ell}(D_{g,1},...,D_{g,1})$ by $Y_{g,1}+\widehat{\Delta}_{g}$. Instead, one could consider an imputation estimator, following Borusyak et al. (2024), Gardner (2021), and Liu et al. (2024). Their papers focus on the case where $D_{1}=0$, but when $D_{1}$ varies we propose to extend their method as follows. For each $d_{1}\in\mathcal{D}^{\text{r}}_{1}$, we regress $Y_{g,t}$ on groups and time fixed effects, on the subsample $\{(g,t):D_{g,t}=d_{1},t<F_{g}\}$. Let $\widehat{\alpha}_{g}$ and $\widehat{\gamma}_{d_{1},t}$ denote the corresponding estimated group and time fixed effects. Then, we impute $Y_{g,F_{g}-1+\ell}(D_{g,1},...,D_{g,1})$ by $\widehat{\alpha}_{g}+\widehat{\gamma}_{D_{g,1},F_{g}-1+\ell}$. Given the asymptotic results on the imputation estimator in Borusyak et al. (2024) when $D_{1}=0$, we anticipate that this estimator would also be asymptotically normal as $G$ tends to infinity, provided that for each $d_{1}$, the number of groups for which $\{(g,t):D_{g,t}=d_{1},t<F_{g}\}$ also goes to infinity.
Finally, in principle we could use other identifying restrictions than parallel trends, e.g. those underlying synthetic controls (Abadie et al., 2010), factor models (Xu, 2017), or synthetic difference-in-differences (Arkhangelsky et al., 2021). Those methods also rely on imputing $Y_{g,t}(D_{1},...,D_{1})$. However, they may be challenging to apply in complex designs, as they require a large number of periods before any treatment change, and a large number of control groups for each set of switchers with a specific value of $D_{1}$.
@@ 3.2.2 Link with effects of specific lags
$\omega^{\ell}_{k}=\frac{S(D_{F-1+\ell-k}-D_{1})}{\text{AVSQ}^{D}_{\ell}}.$
The following result shows that $\text{AVSQ}^{n}_{\ell}$ is the expectation of a weighted average of $(\text{SL}^{\ell}_{k})_{k=0,...,\ell-1}$, with positive weights.
@@
$\text{AVSQ}^{n}_{\ell}=E\left[\sum_{k=0}^{\ell-1}\omega^{\ell}_{k}\text{SL}^{\ell}_{k}\,\big{|}\,\bm{D}\in\mathcal{D}^{\ell}\right].$
@@
This decomposition simplifies much in designs where groups’ treatment can change at most once ($D_{g,t}=D_{g,F_{g}}$ for all $t\geq F_{g}$). Then, $w_{\ell,k}=1/\ell$ for all $k=0,...,\ell-1$. As a result, $\text{AVSQ}^{n}_{1}$ is an effect of the current treatment on the outcome, $\text{AVSQ}^{n}_{2}$ is a weighted average of the effect of the current treatment and of the first treatment lag on the outcome with weights 1/2, and so on. When groups’ treatment can change more than once, estimating $k\mapsto\Omega^{\ell}_{k}:=E[\omega^{\ell}_{k}]$ helps documenting which lags contribute the most to $\text{AVSQ}^{n}_{\ell}$.
@@ 3.2.3 Identification and estimation
$\text{AVSQ}^{n}_{\ell}=\frac{E\left[S(Y_{F-1+\ell}-Y_{F-1}-\Delta_{\ell}(D_{1},F))|\bm{D}\in\mathcal{D}^{\ell}\right]}{E\left[S\sum_{k=0}^{\ell-1}\left(D_{F+k}-D_{1}\right)|\bm{D}\in\mathcal{D}^{\ell}\right]}.$
@@ 3.4 Testing the null that lagged treatments do not affect the outcome
We show here that parameters closely related to the $\text{AVSQ}_{\ell}$ can be used to test the null that there are no dynamic effects, namely:
@@ 3.4.1 Tests when some switchers eventually revert to their initial treatment value
In designs with a binary treatment and where some groups leave the treatment after having been previously treated, Liu et al. (2024) propose a test of Assumption 6 under the standard parallel-trends condition (Assumption 4), which amounts to estimating the average treatment effect across previously treated groups, at time periods where those groups have left the treatment. Under Assumption 6, this average treatment effect should be equal to zero. Their test is implemented by the fect Stata (Liu et al., 2022b) and R (Liu et al., 2022a) commands.
@@
$\text{AVSQ}^{o}_{\ell}:=E\left[S\left(Y_{F+\ell}-Y_{F+\ell}(D_{1},...,D_{1})\right)|D_{F-1+\ell}=D_{1},\bm{D}\in\mathcal{D}^{\ell}\right],$
for all $\ell\geq 2$ for which $P(D_{F-1+\ell}=D_{1},\bm{D}\in\mathcal{D}^{\ell})>0$ (note that $\text{AVSQ}^{o}_{1}$ is undefined since by construction $D_{F}\neq D_{1}$). Under Assumption 6, $Y_{F+\ell}=Y_{F+\ell}(D_{F-1+\ell})$, and thus,
@@ 3.4.2 Tests without switchers reverting to their initial treatment value
but that effect may still depend on $\bm{D}$ in an unrestricted way, if $P(D_{F-1+\ell}=D_{1}|F-1+\ell\leq\overline{T})=0$. Therefore, no-anticipation, parallel trends and Assumption 6 do not have a testable implication.
@@
$\text{AVSQ}^{\text{bal}}_{\ell}:=E[S(Y_{F-1+\ell}-Y_{F-1+\ell}(D_{1},...,D_{1}))|D_{F}=...=D_{F+L},\bm{D}\in\mathcal{D}^{L}],$
@@ 4 Imposing additional restrictions
However, this approach also has limitations. The cost-benefit analysis considered above allows one to compare the actual policy with the status-quo, but cannot evaluate alternative policies, or determine the optimal policy. Also, depending on the design, identification may be limited. When $\ell$ increases, we may be able to find control groups only for very few switchers, and there may also be applications where the no-crossing condition is violated for many switchers. As a result, the subpopulation satisfying $\bm{D}\in\mathcal{D}^{\ell}$ may be small, in particular as $\ell$ increases.
In the rest of this section, we address these issues by considering the following additional restriction:
@@
(Random coefficients distributed-lag linear model) There exist a random vector $B:=(\beta_{0},\dots,\beta_{L})^{\prime}$ such that
$Y_{t}(d_{1},...,d_{t})=Y_{t}(\bm{0}_{t})+\sum_{\ell=0}^{L}\beta_{\ell}d_{t-\ell},$
where L<T1L<T-1 is known. Moreover, $E[\|B\|^{2}]<\infty$.
This assumption is restrictive. It assumes that the analyst knows the number of lags up to which past treatments still affect the current outcome. It also assumes that the effects of the current treatment and its lags are linear, and that they do not complement or substitute each other. Finally, it further assumes that treatment effects do not vary over time. This last condition is also strong, but imposing it seems unavoidable if one wants to use past policy changes to do an ex-ante evaluation of future alternative policies: with time-varying effects, the effects of past policy changes cannot be used to infer the effect of future policies. Note also that Assumption 7 does not impose any restriction on the dependence between the random coefficients $B$ and the treatment path $\bm{D}$.
We first study below what the usual distributed-lag TWFE regressions identify under no anticipation, parallel trends and Assumption 7. Because of the possible correlation between $B$ and $\bm{D}$, such regressions may fail to identify a causal effect. We then consider how to identify expectations of $B$.
@@ 4.1 Distributed-lag two-way fixed effects regressions
In general designs, a commonly-used estimator of dynamic treatment effects, for instance discussed in Equation (5.2.6) of Angrist and Pischke (2009), is the distributed-lag TWFE estimator. For $\ell\in\{0,...,L\}$, let $\widehat{\beta}_{\ell}$ denote the coefficient on $D_{g,t-\ell}$ in a regression of $Y_{g,t}$ on group and period FEs and $\left(D_{g,t-\ell}\right)_{\ell\in\{0,...,L\}}$, in the subsample such that $t\geq L+1$:
$Y_{g,t}=\sum_{g^{\prime}=1}^{G}\widehat{\alpha}_{g^{\prime}}1\{g=g^{\prime}\}+\sum_{t^{\prime}=1}^{T}\widehat{\gamma}_{t^{\prime}}1\{t=t^{\prime}\}+\sum_{\ell=0}^{L}\widehat{\beta}_{\ell}D_{g,t-\ell}+\hat{\epsilon}_{g,t}.$ (11)
@@
Before stating this result, we introduce additional notation. For g=1,,Gg=1,...,G, let $\eta^{\ell}_{g,t}$ denote the sample residual of cell $(g,t)$ ($t\geq L+1$) in the regression of $D_{g,t-\ell}$ on group and time fixed effects and $(D_{g,t-\ell^{\prime}})_{\ell^{\prime}\neq\ell}$. Then, for any $(\ell,\ell^{\prime})\in\{0,...,L\}^{2}$, let
$W^{\ell,\ell^{\prime}}_{g}=\frac{\sum_{t\geq L+1}\eta^{\ell}_{g,t}D_{g,t-\ell^{\prime}}}{(1/G)\sum_{g^{\prime}=1}^{G}\sum_{t\geq L+1}\eta^{\ell}_{g^{\prime}t}D_{g^{\prime},t-\ell}}.$
This random variable depends on $G$ but we omit the dependence to simplify notation. Since the initial sample is i.i.d., the distribution of $W^{\ell,\ell^{\prime}}_{g}$ does not depend on $g$, so we omit this dependence below.
@@
Suppose that Assumptions 1-4 and 7 hold. Then, for all $\ell\in\{0,...,L\}$,
$E\left[\widehat{\beta}_{\ell}\right]=E\bigg{[}W^{\ell,\ell}\beta_{\ell}+\sum_{\begin{subarray}{c}\ell^{\prime}=0\\ \ell^{\prime}\neq\ell\end{subarray}}^{L}\leavevmode\nobreak\ W^{\ell,\ell^{\prime}}\beta_{\ell^{\prime}}\bigg{]}.$ (12)
Moreover, $E[W^{\ell,\ell}]=1$ and $E[W^{\ell,\ell^{\prime}}]=0$ for all $\ell^{\prime}\neq\ell$. Finally, if $\text{Cov}(W^{\ell,\ell^{\prime}},\beta_{\ell^{\prime}})=0$ for all $\ell^{\prime}\in\{0,...,L\}$, $E\left[\widehat{\beta}_{\ell}\right]=E\left[\beta_{\ell}\right]$.
[…]
Importantly, the weights $(W_{g}^{\ell,\ell^{\prime}})_{g=1,...,G}$ can be computed, and this computation can help diagnosing the robustness of the distributed-lag TWFE regression to heterogeneous treatment effects. For instance, a useful diagnostic is to look at the variance of the weights $(W_{g}^{\ell,\ell})_{g=1,...,G}$ around their expectation, equal to one. A large variance indicates that the regression strongly upweights the $\beta_{\ell}$ of some groups and strongly downweights or even weights negatively the $\beta_{\ell}$ of other groups, and if the weights are correlated with groups’ treatment effects, this could bias $\widehat{\beta}_{\ell}$ far away from $E[\beta_{\ell}]$.

### v3

@@ 3.1.3 Placebo effects
$\text{AVSQ}_{-\ell}=E\left[S\times(Y_{F-1-\ell}-Y_{F-1}-\Delta_{-\ell}(D_{1},F))|\bm{D}\in\mathcal{D}^{\ell},F-1-\ell\geq 1\right],$
@@ 3.1.5 Other estimators and identifying strategies
Our estimator of $\text{AVSQ}_{\ell}$ consists in estimating the missing counterfactual outcome $Y_{g,F_{g}-1+\ell}(D_{g,1},...,D_{g,1})$ by $Y_{g,1}+\widehat{\Delta}_{g}$. Instead, one could consider an imputation estimator, following Borusyak et al. (2024), Gardner (2021), and Liu, Wang, and Xu (2024). Their papers focus on the case where $D_{1}=0$, but when $D_{1}$ varies we propose to extend their method as follows. For each $d_{1}\in\mathcal{D}^{\text{r}}_{1}$, we regress $Y_{g,t}$ on groups and time fixed effects, on the subsample $\{(g,t):D_{g,t}=d_{1},t<F_{g}\}$. Let $\widehat{\alpha}_{g}$ and $\widehat{\gamma}_{d_{1},t}$ denote the corresponding estimated group and time fixed effects. Then, we impute $Y_{g,F_{g}-1+\ell}(D_{g,1},...,D_{g,1})$ by $\widehat{\alpha}_{g}+\widehat{\gamma}_{D_{g,1},F_{g}-1+\ell}$. Given the asymptotic results on the imputation estimator in Borusyak et al. (2024) when $D_{1}=0$, we anticipate that this estimator would also be asymptotically normal as $G$ tends to infinity, provided that for each $d_{1}$, the number of groups for which $\{(g,t):D_{g,t}=d_{1},t<F_{g}\}$ also goes to infinity.
Finally, in principle we could use other identifying restrictions than parallel trends, e.g. those underlying synthetic controls (Abadie, Diamond, and Hainmueller, 2010), factor models (Xu, 2017), or synthetic difference-in-differences (Arkhangelsky, Athey, Hirshberg, Imbens, and Wager, 2021). Those methods also rely on imputing $Y_{g,t}(D_{1},...,D_{1})$. However, they may be challenging to apply in complex designs, as they require a large number of periods before any treatment change, and a large number of control groups for each set of switchers with a specific value of $D_{1}$.
@@ 3.2.1 Definition of the parameter
We define placebo effects as $\text{AVSQ}^{n}_{-\ell}:=\text{AVSQ}_{-\ell}/\text{AVSQ}^{D}_{\ell}$ for $\ell\geq 1$. Note that we simply modify the numerator here, since $\text{AVSQ}^{D}_{-\ell}=0$ by construction.
@@ 3.2.2 Link with effects of specific lags
$\omega^{\ell}_{k}=\frac{S\times(D_{F-1+\ell-k}-D_{1})}{\text{AVSQ}^{D}_{\ell}}.$
Because $S\times(D_{F-1+\ell-k}-D_{1})\geq 0$ and $\text{AVSQ}^{D}_{\ell}>0$, $\omega^{\ell}_{k}\geq 0$. Moreover, by definition of $\text{AVSQ}^{D}_{\ell}$, $E\left[\sum_{k=0}^{\ell-1}\omega^{\ell}_{k}|\bm{D}\in\mathcal{D}^{\ell}\right]=1$. Finally, note that
$\sum_{k=0}^{\ell-1}\omega^{\ell}_{k}\text{SL}^{\ell}_{k}=S\times(Y_{F-1+\ell}(D_{1},...,D_{F-1},...,D_{F-1+\ell})-Y_{F-1+\ell}(D_{1},...,D_{1})).$
As a result, we obtain that $\text{AVSQ}^{n}_{\ell}$ is the expectation of a weighted average of $(\text{SL}^{\ell}_{k})_{k=0,...,\ell-1}$, with positive weights:
@@
$\text{AVSQ}^{n}_{\ell}=E\left[\sum_{k=0}^{\ell-1}\omega^{\ell}_{k}\text{SL}^{\ell}_{k}\,\big|\,\bm{D}\in\mathcal{D}^{\ell}\right].$ (8)
@@
The decomposition (8) simplifies much in designs where groups’ treatment can change at most once ($D_{g,t}=D_{g,F_{g}}$ for all $t\geq F_{g}$). Then, $\omega^{\ell}_{k}=1/\ell$ for all $k=0,...,\ell-1$. As a result, $\text{AVSQ}^{n}_{1}$ is an effect of the current treatment on the outcome, $\text{AVSQ}^{n}_{2}$ is a weighted average of the effect of the current treatment and of the first treatment lag on the outcome with weights 1/2, and so on. When groups’ treatment can change more than once, estimating $k\mapsto\Omega^{\ell}_{k}:=E[\omega^{\ell}_{k}|\bm{D}\in\mathcal{D}^{\ell}]$ helps documenting which lags contribute the most to $\text{AVSQ}^{n}_{\ell}$.
The shape of $\ell\mapsto\text{AVSQ}^{n}_{\ell}$ may provide insights as to whether the outcome is more affected by recent or by old treatments. To see this, suppose that $\ell\mapsto\Omega^{\ell}_{k}$ is decreasing for all $\ell>k$ and assume a linear model $Y_{t}(d_{1},...,d_{t})=\mu_{t}+\sum_{k=0}^{t}\delta_{k}d_{t-k}$, in which case $\text{SL}^{\ell}_{k}=\delta_{k}$ for all $k\in\{0,...,\ell-1\}$ and $\ell\geq 1$ for which $P(\bm{D}\in\mathcal{D}^{\ell})>0$. Then, $\text{AVSQ}^{n}_{\ell}=\sum_{k=0}^{\ell-1}\Omega_{k}^{\ell}\delta_{k}$. In this case, one can show that if $k\mapsto\delta_{k}$ is decreasing, $\ell\mapsto\text{AVSQ}^{n}_{\ell}$ is decreasing as well. Hence, a decreasing $\ell\mapsto\text{AVSQ}^{n}_{\ell}$ provides suggestive evidence that potential outcomes are more affected by recent than by past treatments.
@@ 3.2.3 Identification and estimation
$\text{AVSQ}^{n}_{\ell}=\frac{E\left[S\times(Y_{F-1+\ell}-Y_{F-1}-\Delta_{\ell}(D_{1},F))|\bm{D}\in\mathcal{D}^{\ell}\right]}{E\left[S\sum_{k=0}^{\ell-1}\left(D_{F+k}-D_{1}\right)|\bm{D}\in\mathcal{D}^{\ell}\right]}.$
@@ 3.4 Testing the null that lagged treatments do not affect the outcome
While allowing for dynamic effects is appealing, doing so reduces the sample that can be used in the estimation, which may come with high costs in terms of external validity and statistical precision (de Chaisemartin, D’Haultfœuille, Pasquier, Sow, and Vazquez-Bare, 2022). We show here that parameters closely related to the $\text{AVSQ}_{\ell}$ can be used to test the null that there are no dynamic effects, namely:
@@
Not rejecting those tests may suggest we could use estimators ruling out dynamic effects (see, e.g., de Chaisemartin et al., 2022), which may lead to external-validity and precision gains with respect to the estimators discussed in this paper.
@@ 3.4.1 Tests when some switchers eventually revert to their initial treatment value
In designs with a binary treatment and where some groups leave the treatment after having been previously treated, Liu et al. (2024) propose a test of Assumption 6 under the standard parallel-trends condition (Assumption 4), which amounts to estimating the average treatment effect across previously treated groups, at time periods where those groups have left the treatment. Under Assumption 6, this average treatment effect should be equal to zero. Their test is implemented by the fect Stata (Liu, Wang, Xu, Liu, and Liu, 2022b) and R (Liu, Wang, Xu, Liu, and Liu, 2022a) commands.
@@
$\text{AVSQ}^{o}_{\ell}:=E\left[S\left(Y_{F-1+\ell}-Y_{F-1+\ell}(D_{1},...,D_{1})\right)|D_{F-1+\ell}=D_{1},\bm{D}\in\mathcal{D}^{\ell}\right],$
for all $\ell\geq 2$ for which $P(D_{F-1+\ell}=D_{1},\bm{D}\in\mathcal{D}^{\ell})>0$ (note that $\text{AVSQ}^{o}_{1}$ is undefined since by construction $D_{F}\neq D_{1}$). Under Assumption 6, $Y_{F-1+\ell}=Y_{F-1+\ell}(D_{F-1+\ell})$, and thus,
@@ 3.4.2 Tests without switchers reverting to their initial treatment value
but that effect may still depend on $\bm{D}$ in an unrestricted way, if $P(D_{F-1+\ell}=D_{1}|F-1+\ell\leq\overline{T})=0$. Therefore, no-anticipation, parallel trends and Assumption 6 do not have any testable implication.
@@
$\text{AVSQ}^{\text{bal}}_{\ell}:=E[S\times(Y_{F-1+\ell}-Y_{F-1+\ell}(D_{1},...,D_{1}))|D_{F}=...=D_{F+L},\bm{D}\in\mathcal{D}^{L}],$
@@ 4 Imposing additional restrictions
However, this approach also has limitations. While normalized event-study effects estimate weighted averages of marginal effects of the current and lagged treatments on the outcome, they cannot be used to separately estimate those marginal effects. Moreover, the cost-benefit analysis considered above allows one to compare the actual policy with the status-quo, but cannot evaluate alternative policies, or determine the optimal policy.
We address these issues by considering the following additional restriction:
@@
(Random coefficients distributed-lag linear model) There exist a random vector $B:=(\beta_{0},\dots,\beta_{K})^{\prime}$ such that
$Y_{t}(d_{1},...,d_{t})=Y_{t}(\bm{0}_{t})+\sum_{k=0}^{\min(K,t-1)}\beta_{k}d_{t-k},$
where K<T1K<T-1 is known. Moreover, $E[\|B\|^{2}]<\infty$.
Assumption 7 allows for heterogeneous treatment effects across groups, and it does not impose any restriction on the dependence between the random coefficients $B$ and the treatment path $\bm{D}$. On the other hand, Assumption 7 also imposes some restrictions. First, it assumes that the effects of the current treatment and its lags do not vary over time. This condition is strong, but imposing it seems unavoidable if one wants to use past policy changes to do an ex-ante evaluation of future policies: with time-varying effects, the effects of past policy changes cannot be used to infer the effect of future policies. Second, it assumes that the analyst knows the number of lags up to which past treatments still affect the current outcome. Third, it assumes that the effects of the current treatment and its lags are linear, and that they do not complement or substitute each other.
We first study below what the usual distributed-lag TWFE regressions identify under no anticipation, parallel trends and Assumption 7. Because of the possible correlation between $B$ and $\bm{D}$, such regressions may fail to identify a causal effect. We then consider how to identify expectations of $B$. Finally, we consider relaxations of some of the restrictions imposed by Assumption 7.
@@ 4.1 Distributed-lag two-way fixed effects regressions
In general designs, a commonly-used estimator of dynamic treatment effects, for instance discussed in Equation (5.2.6) of Angrist and Pischke (2009), is the distributed-lag TWFE estimator. For $k\in\{0,...,K\}$, let $\widehat{\beta}_{k}$ denote the coefficient on $D_{g,t-k}$ in a regression of $Y_{g,t}$ on group and period FEs and $\left(D_{g,t-k}\right)_{k\in\{0,...,K\}}$, in the subsample such that $t\geq K+1$:
$Y_{g,t}=\sum_{g^{\prime}=1}^{G}\widehat{\alpha}_{g^{\prime}}1\{g=g^{\prime}\}+\sum_{t^{\prime}=1}^{T}\widehat{\gamma}_{t^{\prime}}1\{t=t^{\prime}\}+\sum_{k=0}^{K}\widehat{\beta}_{k}D_{g,t-k}+\hat{\epsilon}_{g,t}.$ (12)
@@
Before stating this result, we introduce additional notation. For g=1,,Gg=1,...,G, let $\eta^{k}_{g,t}$ denote the sample residual of cell $(g,t)$ ($t\geq K+1$) in the regression of $D_{g,t-k}$ on group and time fixed effects and $(D_{g,t-k^{\prime}})_{k^{\prime}\neq k}$. Then, for any $(k,k^{\prime})\in\{0,...,K\}^{2}$, let
$W^{k,k^{\prime}}_{g}=\frac{\sum_{t\geq K+1}\eta^{k}_{g,t}D_{g,t-k^{\prime}}}{(1/G)\sum_{g^{\prime}=1}^{G}\sum_{t\geq K+1}\eta^{k}_{g^{\prime}t}D_{g^{\prime},t-k}}.$
This random variable depends on $G$ but we omit the dependence to simplify notation. Since the initial sample is i.i.d., the distribution of $W^{k,k^{\prime}}_{g}$ does not depend on $g$, so we omit this dependence below.
@@
Suppose that Assumptions 1-4 and 7 hold. Then, for all $k\in\{0,...,K\}$,
$E\left[\widehat{\beta}_{k}\right]=E\bigg[W^{k,k}\beta_{k}+\sum_{\begin{subarray}{c}k^{\prime}=0\\ k^{\prime}\neq k\end{subarray}}^{K}\penalty 10000\ W^{k,k^{\prime}}\beta_{k^{\prime}}\bigg].$ (13)
Moreover, $E[W^{k,k}]=1$ and $E[W^{k,k^{\prime}}]=0$ for all $k^{\prime}\neq k$. Finally, if $\text{Cov}(W^{k,k^{\prime}},\beta_{k^{\prime}})=0$ for all $k^{\prime}\in\{0,...,K\}$, $E\left[\widehat{\beta}_{k}\right]=E\left[\beta_{k}\right]$.
The proof of Theorem 3 is very similar to that of Theorem 2 in de Chaisemartin and D’Haultfœuille (2023), but we include it in the appendix for completeness. It essentially extends Proposition 3 in Sun and Abraham (2021) to non-binary or non-staggered designs. Theorem 3 shows that even under Assumption 7, which is restrictive, the distributed-lag TWFE estimator $\widehat{\beta}_{k}$ does not identify $E\left[\beta_{k}\right]$ in general. Rather, it estimates the sum of $K+1$ terms. The first term is a weighted average of $\beta_{k}$, where the weight variable has expectation one but may be negative. Then, if treatment effects are heterogeneous between groups, we may not have $E[W^{k,k}\beta_{k}]=E[\beta_{k}]$. If $P(W^{k}<0)>0$, we may even have a sign reversal, namely $P(\beta_{k}>0)=1$ yet $E\left[\widehat{\beta}_{k}\right]<0$. Now, (13) also includes a sum of $K$ terms. These terms include $\beta_{k^{\prime}}$, multiplied by another “weight” $W^{k,k^{\prime}}$, which has expectation 0. Thus, $\widehat{\beta}_{k}$, which is supposed to estimate the effect of the $k$th treatment lag, is actually contaminated by the effects of other treatment lags. If $V(\beta_{k^{\prime}})=0$, these contamination terms disappear as the expectation of the contamination weights is equal to zero. But if $V(\beta_{k^{\prime}})>0$, the contamination terms may differ from zero and can also lead to sign reversal.
Importantly, the weights $(W_{g}^{k,k^{\prime}})_{g=1,...,G}$ can be computed, and this computation can help diagnosing the robustness of the distributed-lag TWFE regression to heterogeneous treatment effects. For instance, a useful diagnostic is to look at the variance of the weights $(W_{g}^{k,k})_{g=1,...,G}$ around their expectation, equal to one. A large variance indicates that the regression strongly upweights the $\beta_{k}$ of some groups and strongly downweights or even weights negatively the $\beta_{k}$ of other groups, and if the weights are correlated with groups’ treatment effects, this could bias $\widehat{\beta}_{k}$ far away from $E[\beta_{k}]$.